### v1

Abstract
Accurate detection of Plasmodium falciparum in Giemsa-stained blood smears is an essential component of reliable malaria diagnosis, especially in developing countries. Deep learning-based object detection methods have demonstrated strong potential for automated Malaria diagnosis, but their adoption is limited by the scarcity of datasets with detailed instance-level annotations. In this work, we present an enhanced version of the publicly available NIH malaria dataset, with detailed bounding box annotations in COCO format to support object detection training. We validated the revised annotations by training a Faster R-CNN model to detect infected and non-infected red blood cells, as well as white blood cells. Cross-validation on the original dataset yielded F1 scores of up to 0.88 for infected cell detection. These results underscore the importance of annotation volume and consistency, and demonstrate that automated annotation refinement combined with targeted manual correction can produce training data of sufficient quality for robust detection performance. The updated annotations set is publicly available via GitHub:
https://github.com/MIRA-Vision-Microscopy/malaria-thin-smear-coco.
doi:
10.59275/j.melba.2024-AAAAkeywords:
Malaria, Plasmodium Falciparum, Thin Blood Smear, NIH, COCO1 Background
Malaria is a tropical disease caused by protozoan parasites of the genus Plasmodium, which infect red blood cells and are primarily transmitted through the bites of female Anopheles mosquitoes. In humans, the disease is mainly associated with four species: P. falciparum, P. vivax, P. malariae, and P. ovale. In recent years, malaria has also been increasingly transmitted by a fifth species, i. e., P. knowlesi. Among these five species, P. falciparum and P. vivax are the most prevalent, and P. falciparum is responsible for the majority of malaria-related deaths (World Health Organization, 2024).
Most malaria infections are reported in tropical and subtropical regions, affecting populations in low-income countries with limited access to healthcare. Although modern treatments can effectively cure malaria, early diagnosis remains critical and delays in detection are a major contributing factor to malaria-related mortality (Sultani et al., 2022). For parasitological diagnosis of malaria, microscopic examination of thick and thin blood smear images is routinely performed. In addition to identifying the Plasmodium species, light microscopy allows parasite quantification and monitoring therapy response. Therefore, it is often preferred over molecular testing (World Health Organization, 2024). Nonetheless, the parasitological assessment of blood smear images requires a high level of expertise, and trained personnel might be scarce in low-resource countries or rural areas (Poostchi et al., 2018).
Recently, machine learning-based approaches for analyzing digitized blood smear images have demonstrated promising results in parasitemia quantification (Poostchi et al., 2018). However, these methods typically rely on large, well-annotated datasets for effective training, making publicly available resources particularly valuable. Most existing work focuses on classifying individual cell patches as infected or non-infected (Kassim et al., 2020), which requires the prior extraction of single-cell crops. This step can be challenging in densely populated blood smear images and limits the applicability of such approaches in real-world diagnostic workflows, where direct localization and accurate quantification of infected cells are essential. In contrast to patch-based classification approaches, object detection architectures require datasets with detailed instance-level annotations, typically in the form of labeled bounding boxes. However, acquiring such detailed annotations is labor-intensive and time-consuming, which limits their availability. The NIH dataset, comprising 965 images, is one of the largest publicly available resources for P. falciparum detection. However, only of these images include detailed polygon-based annotations, while the remaining are limited to point annotations marking cell centers. This sparsity limits their suitability for training deep learning-based object detection models, which typically require bounding box annotations.
In this work, we present a revised version of the NIH dataset with enhanced annotations. Using the Cellpose framework (Pachitariu and Stringer, 2022) and manual label correction, we converted the original point annotations into bounding box labels, which are better suited for object detection. To validate the quality of the revised dataset, we trained a Faster R-CNN (Ren et al., 2015) for parasite detection, achieving an F1 score of up to for infected cell identification. The updated annotation set is publicly available via GitHub: https://github.com/MIRA-Vision-Microscopy/malaria-thin-smear-coco
2 Methods
For our experiments, we generated new bounding box annotations for the NIH dataset, which contains Giemsa-stained, thin blood smear images of P. falciparum. We conducted a technical validation of these annotations by training a deep learning-based object detector to identify three cell types: non-infected red blood cells, infected red blood cells, and white blood cells.
2.1 Data Details
The NIH dataset (Kassim et al., 2020) is a thin-smear malaria image dataset acquired at Chittagong Medical College Hospital in Bangladesh and published by the National Library of Medicine, National Institutes of Health (NIH), Bethesda, MD, USA. It comprises Giemsa-stained, thin blood smear images from 193 patients (148 infected and 45 uninfected), with five images per patient. Each image was captured using a microscope-mounted smartphone camera at a resolution of (widthheight) pixels. Annotations cover three classes: non-infected red blood cells, infected red blood cells, and white blood cells. Of the total images, include detailed polygon-based annotations, while the remaining provide only point annotations marking cell centers. Table 1 summarizes these subsets, hereafter referred to as NIH and NIH, respectively. Figures 1(a) and 1(b) show example regions of interest with contour and point annotations, corresponding to the NIH and NIH subsets.
| NIH | NIH | MIRA | |
| patients | |||
| no. of images | |||
| annotations | contours | points | boxes |
| no. of annotations | |||
| non-infected | |||
| infected | |||
| white blood cell | |||
| ambiguous | - | - |
2.2 Annotation Revision
To enable the use of the NIH dataset for training object detection models, we converted the point annotations into detailed bounding-box annotations. For this, we first detected cell instances using Cellpose 2 (Pachitariu and Stringer, 2022), an open-source framework designed for robust, generalizable segmentation. Trained with a diverse dataset of more than cells, Cellpose offers strong performance across a wide range of cell types and imaging modalities, making it well suited for segmenting Giemsa-stained blood smear images.
Following cell instance segmentation, we assigned labels to detected cells by overlaying the original point annotations. If a point annotation fell within a predicted bounding box, that box was assigned the corresponding cell class. However, Cellpose occasionally detected cells, which were not annotated in the original dataset. These were often partially visible cells at the edge of the field of view. In the updated annotation set, these detections were labeled as ambiguous. Figure 2 shows an example with ambiguous cells at the border of the field of view. Overall, the updated annotations comprise ambiguous cells, which makes up around of the original NIH subset.
Due to its reliance on an average cell size, Cellpose sometimes fragmented larger cells and particularly white blood cells into multiple instances. To address this, we manually reviewed and merged these fragmented detections. Additionally, Cellpose occasionally misclassified artifacts or blood platelets as cells. These false positives were also removed during manual post-processing. Figure 1(c) shows a representative region of interest after bounding box detection, with ambiguous cells highlighted in orange, and the last column of Table 1 summarizes the number of cell instances after this annotation revision.
3 Technical Validation
To validate the revised annotations, we trained a Faster R-CNN model (Ren et al., 2015) to detect three cell classes: non-infected red blood cells, infected red blood cells, and white blood cells. We conducted cross-validation experiments by training the model on either the NIH or the revised MIRA subset and evaluating its detection performance on the other, respectively.
3.1 Implementation Details
We employed a Faster R-CNN model (Ren et al., 2015) with a ResNet34 (He et al., 2016) backbone, pretrained on ImageNet (Russakovsky et al., 2015). The datasets were split into for training and for validation. On the NIH dataset, the model was trained for epochs using a cosine annealing learning rate schedule with linear warm-up over the first 50 epochs and a maximum learning rate of . For the MIRA dataset, the training time was lowered to epochs, to match the almost five-fold size of the data subset. For optimization, the Adam optimizer and standard Faster R-CNN loss functions were used. Training patches of pixels were sampled from the original pixel images. This resolution was chosen to match the 4:3 aspect ratio typical of microscopy images, while ensuring that each patch contained a sufficient number of cells for effective training. The patches were then downscaled by a factor of 2 to a final size of pixels, enabling a batch size of without exceeding memory constraints. To address class imbalance, we applied a custom patch sampling strategy that over-sampled regions containing underrepresented classes, such as white blood cells. Model performance was monitored using the mean average precision (mAP) on the validation set, and the final model was selected based on the best validation mAP.
For inference on the full-resolution pixel images, we used the SAHI framework (Akyon et al., 2021, 2022), which performs sliding-window predictions and applies non-maximum suppression (NMS) to eliminate duplicate detections across overlapping patches. As a post-processing step, we removed all predicted bounding boxes with an area smaller than pixels or larger than pixels. These thresholds were determined based on the minimum and maximum annotation sizes observed in the original NIH dataset.
3.2 Evaluation
For evaluation, we computed class-wise F1 scores from the instance-level confusion matrices. Cells that were detected by Cellpose but not labeled by human annotators (i. e., ambiguous cells) were excluded from the evaluation. Annotated cells that were not detected by the model were considered false negatives due to detection failure (), while model predictions that were not annotated and not labeled as ambiguous were considered false positives due to detection failure (). The class-wise F1 score for class was computed as:
3.3 Results
Figure 3(a) presents the confusion matrix of the Faster R-CNN model trained on the NIH subset and evaluated on the MIRA subset, and vice versa. Results are displayed as row-normalized percentages along with absolute cell counts.
Overall, the model performs better when trained on the MIRA and evaluated on the NIH subset than the other way round, reflected by a lower proportion of off-diagonal entries in the confusion matrix. When training on NIH and testing on MIRA, a comparably high ratio () of infected cells was misclassified as non-infected, indicating reduced recall for malaria detection. Furthermore, of the cells annotated as ambiguous were not detected by the model. Closer inspection of these cells revealed that ambiguous cells were often located near the field-of-view borders, where annotations were inconsistently applied. Specifically, these border cells were frequently unannotated in both, the NIH and NIH dataset. This suggests a possible labeling bias, which is further illustrated in Fig. 4, where white arrows indicate unlabeled yet clearly visible cells.
Table 2 summarizes the detection performance, reported as precision, recall, and F1 scores computed from the confusion matrices according to Eqs. 1, 2 and 3. For each dataset, training was repeated with three different random seeds, and we report the average performance as mean standard deviation ().
| NIH MIRA | MIRA NIH | |
| Precision | ||
| non-infected cells | p m 0.01 | p m 0.00 |
| infected cells | p m 0.01 | p m 0.01 |
| white blood cells | p m 0.04 | p m 0.03 |
| Recall | ||
| non-infected cells | p m 0.01 | p m 0.00 |
| infected cells | p m 0.01 | p m 0.01 |
| white blood cells | p m 0.03 | p m 0.00 |
| F1 score | ||
| non-infected cells | p m 0.01 | p m 0.00 |
| infected cells | p m 0.01 | p m 0.00 |
| white blood cells | p m 0.04 | p m 0.02 |
The results demonstrate high performance for the detection of non-infected red blood cells and white blood cells, with F1 scores above . Infected cells were detected with an average F1 score of , when training on NIH and when training on MIRA, indicating good but comparatively lower performance. The repeated training runs demonstrate low variability, indicated by a low standard deviation of performance results.
The performance metrics again demonstrate a superior performance of the model trained on MIRA. This especially holds for the recall of infected cells, with average values of for training on NIH and for training on MIRA. This observation could be attributed to discrepancies in labeling consistency, but also to the higher volume of annotated instances ( vs. ), which provides more diverse training examples to the model.
4 Discussion and Summary
This study presents a revised version of the NIH malaria dataset with instance-level annotations in COCO format, facilitating the development of deep learning-based object detection models for the automatic detection of infected cells. We validated these annotations by training a Faster R-CNN to detect infected and non-infected red blood cells, as well as white blood cells, achieving an F1 score of up to for the detection of infected cells. For trained microscopists, the World Health Organization (WHO) guidelines recommend a minimum recall of infected malaria samples of (World Health Organization, 2009), which our system achieves on a cellular level. Therefore, the system meets the minimum competency level required in a diagnostic setting. Nevertheless, our analysis of ambiguous cells revealed inconsistencies in the original annotations, where especially at the image borders cells were not labeled by the pathologists. However, it is difficult to tell whether these cells were simply overlooked or not labeled on purpose as a reliable malaria diagnosis might not be possible on partially visible cells. This raises broader concerns about ground truth quality in biomedical datasets likely caused by a trade-off of labeling precision and time investment. To the best of our knowledge, the original dataset was annotated by a single expert, which can introduce a considerable labeling bias. Future work could address this with additional manual annotation rounds with consensus labeling by multiple experts or introducing a separate class for partially visible cells. For evaluating the performance of machine learning models, we recommend excluding these cells from evaluation.
Despite the challenges associated with partially labeled data, our results demonstrate that annotation conversion via existing tools such as Cellpose, followed by targeted manual curation, can yield training data of sufficient quality to support robust model performance. This finding is particularly relevant for resource-constrained settings where detailed annotations are expensive or infeasible.
In addition to annotation consistency, we also observed differences in model performance between the two subsets of the NIH dataset, likely driven by the varying number of annotated instances available for training. This highlights the importance of dataset size and diversity for learning subtle morphological features, such as the presence of ring-stage parasites. Furthermore, our initial dataset assessment demonstrated a high class imbalance of healthy and infected cells. We compensated for this to some extent by employing a customized patch sampling strategy, but in future work dedicated augmentation strategies or class-balanced loss functions could be integrated.
Overall, our work contributes an enhanced dataset and a robust pipeline for parasite detection in microscopy, supporting further research into automated malaria diagnosis.
The authors do not have any conflicts of interest to declare.
The authors acknowledge the U.S. National Library of Medicine for making the thin blood smear dataset used in this work publicly available. The dataset is provided under a license that permits redistribution and modification, with appropriate attribution. The authors gratefully acknowledge the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU). The hardware is funded by the German Research Foundation (DFG).
The work follows appropriate ethical standards in conducting research and writing the manuscript, in accordance with all applicable laws and regulations regarding the treatment of human subjects. The dataset used in this study was publicly released by Kassim et al. (2020), who obtained the necessary ethical approvals as documented in the original publication. No additional ethical approval was required.
The dataset (Kassim et al., 2020) used in this study was developed and funded by the U.S. National Library of Medicine (NML), part of the National Institutes of Health (NIH), and is publicly available for commercial and non-commercial use. Use of this dataset is governed by a license that requires proper attribution. We acknowledge the source of the data as follows: “Courtesy of the U.S. National Library of Medicine.” The dataset and associated information are available at: https://lhncbc.nlm.nih.gov/publication/pub9932. Please cite the dataset as described by Kassim et al. (2020). The updated annotation set with the modifications described in Section 2.2 is publicly available via GitHub: https://github.com/MIRA-Vision-Microscopy/malaria-thin-smear-coco.